# Antithetic Noise in Diffusion Models

### v1

@@ 3.2 Explanatory experiments: temporal correlations & symmetry conjecture
Throughout the transition from noises to samples, the PN correlation of $\epsilon_{\theta}^{(t)}$—indicated by the orange, blue, and green dashed lines in Figure 3—starts at –1, stays strongly negative, and only climbs slightly in the final steps. This nearly $-1$ correlation is remarkable, as it suggests that the score network $\epsilon_{\theta}$ learns an approximately affine antisymmetric function. To explain, we first state the following results which shows $-1$ correlation is equivalent to affine antisymmetric, with proof in Appendix A.1.
Figure 3: Correlation of $x_{t}$ (solid) and $\epsilon_{\theta}^{(t)}$ (dash) between antithetic (PN) pairs. Step 50 is the initial noise and Step 0 is the generated image. Shaded bands show ±$\pm 1$ std. dev.
@@
Lemma 1 and Figure 3 lead us to the following conjecture:
@@
The conjecture implies that the score network has learned an “almost odd” function up to an additive shift. This is consistent with both theory and observations for large $t$, where $p_{t}(\mathbf{x})$ is close to standard Gaussian, thus the score function is almost linear in $\mathbf{x}$. Nonetheless, our findings illustrate that this conjecture is far more general: it applies to every timestep $t$ and can accommodate genuinely nonlinear odd behaviors (for example, functions like $\sin x$).
@@
To (partly) validate our conjecture, we perform the following experiment. Using a pretrained CIFAR-10 score network with a 50-step DDIM sampler, we picked time steps $t=1,3,\dots,19$ (where smaller $t$ is closer to the image and larger $t$ is closer to pure noise). For each $t$, we sample $\mathbf{x}\sim\mathbb{N}(0,I_{d})$, evaluated the first coordinate of $\epsilon_{\theta}^{(t)}(c\,\mathbf{x})$ as $c$ varied from $-1$ to $1$ (interpolating from $-\mathbf{x}$ to $\mathbf{x}$), and plotted the result in Figure 4. We exclude $t\geq 20$, where the curve is very close to a straight line.
htbp! Refer to caption
@@
Figure 4 confirms that, at every step $t$, the first coordinate of the score network is indeed overall affine antisymmetric. Although the curves display nonlinear oscillations at small $t$, the deviations on either side are approximately mirror images, and for larger $t$ the mapping is almost a straight line. The symmetry center $c_{t}$ usually lies near zero, but not always—for instance, at $t=11$, $c_{t}\approx 0.05$ even though the function spans from $-0.05$ to $0.15$. Additional validation experiments—using different coordinates and alternative datasets—are presented in the Appendix B.2. Despite the distinct function shapes across coordinates, the conjectured symmetry still persists.
@@
Both the conjecture and its empirical support could be of independent interest. They imply that the score network learns a function with strong symmetry, even though that symmetry is not explicitly enforced. The conjecture matches the Gaussian linear score approximation in the high-noise regime, which has proven effective wang2024the . In the intermediate-to-low noise regime, a single Gaussian approximation performs poorly—this is clear in Figure 4 for $t=3,7,11$, where the score is strongly nonlinear—yet our conjectured approximate symmetry still holds. We believe this result provides valuable insights into pretrained score networks, which are commonly perceived as black boxes.
This experiment is mainly intended to explain our negative correlation finding and should be seen as initial evidence for the conjecture. Fully validating or disproving the symmetry conjecture would require much more extensive experiments, which we leave as future work.
3.3 An alternative explanation via the FKG inequality
[…]
To generalize the result to higher dimension, we define a function $f:\mathbb{R}^{m}\to\mathbb{R}$ to be partially monotone if for each coordinate $j$, holding all other inputs fixed, the map $t\;\mapsto\;f(x_{1},\dots,x_{j-1},t,x_{j+1},\dots,x_{m})$ is either non-decreasing or non-increasing. Mixed monotonicity is allowed. For example, $f(x,y)=x-y$ is non-decreasing in $x$ and non-increasing in $y$, yet still qualifies as partially monotone. We have the following result:
Proposition 1.
Our proof is based on an adaptation of the FKG inequality and is provided in Appendix A.2.
Proposition 1 relies on checking partial-monotonicity. If $S$ is partially monotone (including any linear statistic), then the conditions of Proposition 1 are satisfied by, e.g., Neural Additive Models agarwal2021neural and Deep Lattice Networks you2017deep . Unfortunately, partial-monotonicity is in general hard to verify.
@@ 4 Uncertainty quantification
To assess a diffusion model, one typically draws multiple initial noises, computes the desired statistic for each output, and aggregates the results into point estimates and confidence intervals, which also enables uncertainty estimation from multiple samples beyond a single run. Common statistics of interest include (i) mean pixel intensity for detecting signal leakage lin2024common ; everaert2024exploiting and out-of-distribution detection le2024detecting , (ii) CLIP similarity for measuring semantic alignment in text-to-image models bent2024semantic ; meral2024conform , and (iii) reconstruction error across posterior samples for assessing performance and uncertainty luo2023bayesian ; wu2024unsupervised .
Our goal is to leverage negative correlation to design an estimator with higher accuracy at fixed cost.
Standard Monte Carlo: Formally, consider a diffusion model DM that maps noise to data, and a statistic $S$. The quantity of interest is $\mathbb{E}_{z\sim\mathbb{N}(0,I)}[S(\text{DM}(z))]$. To approximate this expectation, the simplist approach is to draw $N$ independent noise samples $z_{1},\dots,z_{N}\sim\mathbb{N}(0,I)$, calculate $S_{i}=S\bigl{(}\text{DM}(z_{i})\bigr{)}$ for each $i$ and form the standard Monte Carlo (MC) estimator $\hat{\mu}_{N}^{\text{MC}}\coloneq\sum_{i=1}^{N}S_{i}/N$ by taking their average. A $(1-\alpha)$ confidence interval, denoted as $\text{CI}^{\text{MC}}_{N}(1-\alpha)$ is then
$\hat{\mu}_{N}^{\text{MC}}\;\pm\;z_{1-\alpha/2}\,\sqrt{\frac{(\hat{\sigma}_{N}^{\text{MC}})^{2}}{N}},$
where $\hat{\sigma}_{N}^{2}$ is the sample variance of $S_{1},S_{2},\ldots,S_{N}$ and $z_{1-\alpha/2}$ is the $(1-\alpha/2)$-quantile of the standard normal. A formal guarantee of the above construction is given below:
Proposition 2.
Antithetic Monte Carlo: The observed negative correlation motivates an improved estimator. Let $N=2K$ be even, users can generate $K$ pairs of antithetic noise $(z_{1},-z_{1}),\ldots,(z_{K},-z_{K})$. Define $S_{i}^{+}=S(\text{DM}(z_{i}))$, $S_{i}^{-}=S(\text{DM}(-z_{i}))$, and let $\bar{S}_{i}=0.5(S_{i}^{+}+S_{i}^{-})$ be their average. Our Antitietic Monte Carlo (AMC) estimator is $\hat{\mu}_{N}^{\text{AMC}}:=\sum_{i=1}^{K}\bar{S}_{i}/K$, and confidence interval $\text{CI}_{N}^{\text{AMC}}(1-\alpha)$ is
$\vspace{-0.3em}\hat{\mu}_{N}^{\text{AMC}}\;\pm\;z_{1-\alpha/2}\,\sqrt{\frac{2(\hat{\sigma}_{N}^{\text{AMC}})^{2}}{N}},$
where $(\hat{\sigma}_{N}^{\text{AMC}})^{2}$ is the sample variance of $\bar{S}_{1},\ldots,\bar{S}_{K}$.
@@
The idea of using negatively correlated samples has been widely adopted in Monte Carlo methods (craiu2005multiprocess, ) and statistical risk estimation liu2024crossvalidation for improved performance. Generalizing beyond paired correlation, one could generate $N$ negatively correlated Gaussian noises $z_{1},\ldots,z_{N}$ with $\mathrm{Corr}(z_{j},z_{k})=-1/(N-1)$ for $j\neq k$. When $N=2$, this recovers our antithetic sampling.
@@
This principle of variance reduction can be further extended to QMC methods. As an alternative to Monte Carlo, QMC constructs a deterministic set of $N$ samples that have negative correlation and provide a more balanced coverage of the sample space. QMC samples can also be randomized, resulting in RQMC methods. RQMC maintains the marginal distribution of each sample while preserving the low-discrepancy properties of QMC. Importantly, repeated randomizations allow for empirical error estimation and confidence intervals lecuyer2023confidence .
@@
With negative correlation ($\rho<0$), AMC improves the variance by a constant factor of $1+\rho$. In contrast, under regularity conditions nied:1992 , QMC and RQMC can improve the error convergence rate from $O(N^{-1/2})$ to $O(N^{-1+\epsilon})$ for any $\epsilon>0$, albeit $\epsilon$ may absorb a dimension-dependent factor $(\log N)^{d}$. For sufficiently smooth functions, RQMC can even achieve rates as fast as $O(N^{-3/2+\epsilon})$ snetvar ; smoovar . Moreover, the space-filling nature of QMC may also lead to more diverse sample generation.
@@
Although the dimension of the Gaussian noise is higher than the typical regime where QMC is expected to be most effective, RQMC still further shrinks our confidence intervals by several-fold compared with standard MC. This hints that the image generator’s effective dimension is much lower than its ambient one, allowing QMC methods to stay useful. Similar gains arise when QMC methods deliberately exploit low-dimensional structure in practical problems wang2003effective ; xiao:wang:2019 ; liu2023preintegration . Developing ways to identify and leverage this structure more systematically is an appealing avenue for future work.
5 Experiments
We apply the antithetic-noise approach to the two tasks introduced in Section 1: image diversity and uncertainty quantification. Throughout this section, we evaluate our method across three settings: (1) unconditional diffusion models on standard datasets, (2) text-to-image generation using Stable Diffusion, and (3) measurement-guided generation via Diffusion Posterior Sampling (DPS).
We use pretrained models from Hugging Face’s Diffusers library von-platen-etal-2022-diffusers : google/ddpm-church-256, google/ddpm-cifar10-32, and google/ddpm-celebahq-256 for unconditional diffusion; Stable Diffusion v1.5 for text-to-image; and the original repository from chung2023diffusion for guided generation. Experiments were run on eight NVIDIA L40 GPUs. The most intensive setup—Stable Diffusion—takes about five minutes to generate 100 images for a single prompt.
5.1 Improving diversity
Many diffusion tools generate a few images from one prompt to let users select a prefered one. For example, DALL·E 3 via ChatGPT interface outputs two images, Kera.ai outputs three, and Midjourney outputs four. However, these images often look similar and are thus less useful marwood2023diversity ; sadat2024cads .
We show that antithetic noise yields more diverse images compared with independent noise, with no loss in image quality or extra computational cost. We expect our method can be easily integrated into existing diversity optimization techniques to further improve their performance. We quantify diversity by computing pairwise SSIM to measure structural similarity (lower = more diverse) and LPIPS zhang2018unreasonable (higher = more diverse) to capture perceptual distance in deep feature space.
We begin with unconditional image generation using pre-trained diffusion models on standard datasets. We prepare 1,600 PN noise pairs (antithetic) and 1,600 RR noise pairs (independent) as inputs and generate images with a 50-step DDIM sampler.
The diversity metrics are computed for each pair. As shown in Table 2, antithetic noise pairs consistently lead to higher diversity, as indicated by lower SSIM and higher LPIPS scores. Importantly, image quality remains stable for both noise types. For example, on CIFAR10, the mean MUSIQ scores are 16.82 for PN and 16.78 for RR. This suggests that the increased diversity introduced by antithetic noise does not come at the cost of visual quality.
We next evaluate our method using Stable Diffusion with 20-step DDIM. We evaluate the model on a total of 200 prompts selected from the Pick-a-Pic kirstain2023pickapic and DrawBench benchmarks saharia2022photorealistic . For each prompt, we generate 100 image pairs using antithetic and independent initial noises respectively. As in the unconditional setting, antithetic noise consistently results in greater diversity, shown in Table 2, while still preserving semantic alignment with the prompt indicated by CLIP scores of 30.07 (PN) and 30.06 (RR). Notably, the improvement holds for nearly all prompts; for instance, SSIM improved in 199 out of 200 prompts, and we suspect the only remaining case is due to statistical fluctuation.
Metric Unconditional Diffusion Stable Diffusion
CIFAR-10 LSUN-Church CelebA-HQ DrawBench Pick-a-Pic
SSIM (%) 86.1422 34.0056 32.8561 21.511 19.0020
LPIPS (%) 8.2881 4.2029 18.1126 4.8170 5.3613
5.2 Uncertainty quantification
We compute various statistics using MC, AMC, and QMC estimators discussed in Section 4.
We define the relative efficiency of method A with respect to the standard MC baseline as the squared ratio of their confidence interval (CI) lengths, $(\text{CI}_{\text{MC}}/\text{CI}_{\text{A}})^{2}$. Since standard MC produces CIs of length $O(N^{-1/2})$, the relative efficiency quantifies how many more MC samples are required to match the A’s CI length. For each method, we report both the 95% CI width and its relative efficiency.
5.2.1 Estimating pixel-wise statistics
The custom statistics include: pixel value mean (the average intensity across all pixels), perceived brightness (computed via a luminance-weighted combination of RGB channels), contrast (the difference in brightness between the top and bottom halves of the image), and image centroid (the vertical coordinate of the intensity mass center), detailed definitions are in Appendix B.3.
To ensure equal cost, we fix a budget of 3,200 samples for unconditional diffusion model and 200 samples/prompt for Stable Diffusion (across 200 prompts). Therefore, for unconditional diffusion, the MC estimator uses 3,200 independent samples; the AMC estimator uses 1,600 pairs from antithetic noises ($1,600\times 2=3,200$); and QMC uses a Sobol’ point set of size 64 and applies 50 independent randomization ($50\times 64=3,200$). For Stable Diffusion, MC uses 200 independent samples; AMC uses 100 pairs; and QMC uses 8 Sobol’ points and 25 independent randomization. For MC and AMC, CIs are constructed as discussed in Section 4. For QMC, since the number of replicates is relatively small, we use the student-$t$ CI to ensure reliable coverage  lecuyer2023confidence . The Sobol’ sequence uses direction numbers from joe2008constructing , where the maximum number of dimensions is 21,201, limiting our QMC implementation to CIFAR-10 and Stable Diffusion. Details on RQMC CIs are in Appendix B.3.

### v2

@@ 3.2 Explanatory experiments: temporal correlations & symmetry conjecture
Throughout the transition from noises to samples, the PN correlation of $\epsilon_{\theta}^{(t)}$—indicated by the orange, blue, and green dashed lines in Figure 3—starts at –1, stays strongly negative, and only climbs slightly in the final steps. This nearly 1-1 correlation is remarkable, as it suggests that the score network $\epsilon_{\theta}$ learns an approximately affine antisymmetric function. To explain, we first state the following results which shows 1-1 correlation is equivalent to affine antisymmetric, with proof in Appendix A.1.
Figure 3: Correlation of $x_{t}$ (solid) and $\epsilon_{\theta}^{(t)}$ (dash) between antithetic (PN) pairs. Step 50 is the initial noise and Step 0 is the generated image. Shaded bands show $\pm 1$ std. dev.
@@
Lemma 1 and Figure 3 leads us to the following conjecture:
@@
The conjecture implies that the score network has learned an “almost odd” function up to an additive shift. This is consistent with both theory and observations for large $t$, where $p_{t}(\mathbf{x})$ is close to standard Gaussian, thus the score function is almost linear in $\mathbf{x}$. Nonetheless, our conjecture is far more general: it applies to every timestep $t$ and can accommodate genuinely nonlinear odd behaviors (for example, functions like $\sin x$).
@@
To (partly) validate our conjecture, we perform the following experiment. Using a pretrained CIFAR-10 score network with a 50-step DDIM sampler, we picked time steps $t=1,3,\dots,19$ (where smaller $t$ is closer to the image and larger $t$ is closer to pure noise). For each $t$, we sample $\mathbf{x}\sim\mathbb{N}(0,I_{d})$, evaluated the first coordinate of $\epsilon_{\theta}^{(t)}(c\,\mathbf{x})$ as $c$ varied from 1-1 to $1$ (interpolating from $\mathbf{x}$ to $\mathbf{x}$), and plotted the result in Figure 4. We exclude $t\geq 20$, where the curve is very close to a straight line.
@@
Figure 4 confirms that, at every step $t$, the first coordinate of the score network is indeed overall affine antisymmetric. Although the curves display nonlinear oscillations at small $t$, the deviations on either side are approximately mirror images, and for larger $t$ the mapping is almost a straight line. The symmetry center $c_{t}$ usually lies near zero, but not always—for instance, at $t=11$, $c_{t}\approx 0.05$ even though the function spans from 0.05-0.05 to $0.15$. In Appendix B.5, we present further validation experiments using alternative coordinates, datasets, and a quantitative metric called the antisymmetry score. Despite the distinct function shapes across coordinates, the conjectured symmetry still persists.
We further evaluate the antisymmetry conjecture across additional coordinates and datasets. On CIFAR-10 and Church, we assess how closely one-dimensional slices of the network outputs behave like affine-antisymmetric functions. To this end, we introduce the affine antisymmetry score, which quantifies the degree of antisymmetry. Across datasets, the resulting antisymmetry scores are consistently high (mean values above 0.99 and even the lower quantiles still near 0.9). These results provide strong empirical support for the conjecture; full experimental details and plots are given in Appendix B.5.
Finally, we provide theoretical support for our conjecture in Appendix A.2, A.3, and A.4. Appendix A.2 confirms the conjecture in the large-$t$ regime, Appendix A.3 shows how both the density ratio and the score error converge monotonically as $t$ evolves via a Hermite polynomial expansion, and Appendix A.4 demonstrates that all orthogonal symmetries of the data distribution are preserved throughout the forward process. These results help explain the behavior underlying the conjecture. The forward process preserves every orthogonal symmetry of the initial distribution, including coordinate reflections. Thus, in the ideal case where the data distribution at $t=0$ is reflection-symmetric, the density remains even and the score remains odd for all $t$ by Proposition 4. Moreover, even if this symmetry holds only approximately for small $t$, the monotone decay of the density ratio and the score error shows that $p_{t}$ quickly approaches the Gaussian, so the forward dynamics push the score toward an odd function in a controlled manner.
@@
Both the conjecture and its empirical support could be of independent interest. They imply that the score network learns a function with strong symmetry, even though that symmetry is not explicitly enforced. The conjecture matches the Gaussian linear score approximation in the high-noise regime, which has proven effective (Wang and Vastola, 2024). In the intermediate-to-low noise regime, a single Gaussian approximation performs poorly—this is clear in Figure 4 for $t=3,7,11$, where the score is strongly nonlinear—yet our conjectured approximate symmetry still holds.
Besides explaining the source of negative correlation, we believe this result provides new insight into the structure of diffusion models, which are often treated as black-box functions. Leveraging this finding for algorithms and applications is an important direction for future work.
For readers interested in an additional theoretical perspective, Appendix A.6 presents a discussion that links negative correlation to the FKG inequality, and provides a detailed analysis of the DDIM sampler using this framework.
@@ 4 Uncertainty quantification
In uncertainty quantification for a diffusion model DM, the goal is typically to estimate expectations of the form $\mathbb{E}_{z\sim\mathbb{N}(0,I)}\!\left[S\bigl(\text{DM}(z)\bigr)\right],$ where $S$ is a statistic of user’s interest. Our goal is to leverage negative correlation to design an estimator with higher accuracy at fixed computation cost of inference.
Recent studies have examined epistemic uncertainty and bias in diffusion models, including Berry et al. (2024, 2025). Our focus is different: we study aleatoric uncertainty from noise sampling and develop a variance-reduction method.
Standard Monte Carlo: To approximate this expectation, the simplest approach is to draw $N$ independent noises $z_{1},\dots,z_{N}\sim\mathbb{N}(0,I)$, calculate $S_{i}=S\bigl(\text{DM}(z_{i})\bigr)$ for each $i$ and form the standard Monte Carlo (MC) estimator $\hat{\mu}_{N}^{\text{MC}}\coloneq\sum_{i=1}^{N}S_{i}/N$ by taking their average. A $(1-\alpha)$ confidence interval, denoted as $\text{CI}^{\text{MC}}_{N}(1-\alpha)$ is then $\hat{\mu}_{N}^{\text{MC}}\;\pm\;z_{1-\alpha/2}\,\sqrt{(\hat{\sigma}_{N}^{\text{MC}})^{2}/N},$ where $\hat{\sigma}_{N}^{2}$ is the sample variance of $S_{1},S_{2},\ldots,S_{N}$ and $z_{1-\alpha/2}$ is the $(1-\alpha/2)$-quantile of the standard normal. A formal guarantee of the above construction is given in Proposition 5 in Appendix A.5.
Antithetic Monte Carlo: The observed negative correlation motivates an improved estimator. Let $N=2K$ be even, users can generate $K$ pairs of antithetic noise $(z_{1},-z_{1}),\ldots,(z_{K},-z_{K})$. Define $S_{i}^{+}=S(\text{DM}(z_{i}))$, $S_{i}^{-}=S(\text{DM}(-z_{i}))$, and let $\bar{S}_{i}=0.5(S_{i}^{+}+S_{i}^{-})$ be their average. Our Antithetic Monte Carlo (AMC) estimator is $\hat{\mu}_{N}^{\text{AMC}}:=\sum_{i=1}^{K}\bar{S}_{i}/K$, and confidence interval $\text{CI}_{N}^{\text{AMC}}(1-\alpha)$ is $\hat{\mu}_{N}^{\text{AMC}}\;\pm\;z_{1-\alpha/2}\,\sqrt{2(\hat{\sigma}_{N}^{\text{AMC}})^{2}/N},$ where $(\hat{\sigma}_{N}^{\text{AMC}})^{2}$ is the sample variance of $\bar{S}_{1},\ldots,\bar{S}_{K}$.
@@
$K$-antithetic noise: We can generalize the antithetic noise pair to a collection of $K$ noise variables, constructed so that every pair has the same negative correlation 1/(K1)-1/(K-1). One way to generate them is as follows. Draw $K$ independent standard Gaussian vectors $(w_{1},\ldots,w_{K})$ and let $\bar{w}$ be their average. For each $i$, set $z_{i}=\sqrt{K/(K-1)}\left(w_{i}-\bar{w}\right).$ One can directly check that each $z_{i}$ is still marginally standard Gaussian, and $\operatorname{Corr}(z_{i,l},z_{j,l})=-1/(K-1)$ for every $i\neq j$ and every pixel $l$. In particular, when $K=2$, this construction reduces to our usual antithetic noise pair.
@@
The idea of using negatively correlated samples has been widely adopted in Monte Carlo methods (Craiu and Meng, 2005) and statistical risk estimation (Liu et al., 2024) for improved performance. This principle of variance reduction can be further extended to QMC methods. As an alternative to Monte Carlo, QMC constructs a deterministic set of $N$ samples that have negative correlation and provide a more balanced coverage of the sample space. QMC samples can also be randomized, resulting in RQMC methods. RQMC maintains the marginal distribution of each sample while preserving the low-discrepancy properties of QMC. Importantly, repeated randomizations allow for empirical error estimation and confidence intervals (L’Ecuyer et al., 2023).
@@
Under regularity conditions (Niederreiter, 1992), QMC and RQMC can improve the error convergence rate from $O(N^{-1/2})$ to $O(N^{-1+\epsilon})$ for any $\epsilon>0$, albeit $\epsilon$ may absorb a dimension-dependent factor $(\log N)^{d}$. For sufficiently smooth functions, RQMC can achieve rates as fast as $O(N^{-3/2+\epsilon})$ (Owen, 1997a, b). Moreover, the space-filling property of QMC may also promote greater sample diversity.
@@
Although the dimension of the Gaussian noise is higher than the typical regime where QMC is expected to be most effective, RQMC still further shrinks our confidence intervals by several-fold compared with standard MC. This hints that the image generator’s effective dimension is much lower than its ambient one, allowing QMC methods to stay useful. Similar gains arise when QMC methods deliberately exploit low-dimensional structure in practical problems (Wang and Fang, 2003; Xiao and Wang, 2019; Liu and Owen, 2023). Developing ways to identify and leverage this structure more systematically is an appealing avenue for future work.
5 Experiment
Sections 5.1 and 5.2 present two uncertainty quantification applications: estimating pixel-wise statistics and evaluating diffusion inverse problem solvers. The latter considers two popular algorithms across a range of tasks and datasets. For each task, we apply MC, AMC, and, when applicable, QMC estimators as described in Section 4. For each estimator, we report the $95\%$ confidence interval (CI) width and its relative efficiency. The relative efficiency of a new estimator (AMC or QMC) is defined as the squared ratio of the MC CI width to that of the estimator, $\left(\text{CI}_{\text{MC}}/\text{CI}_{\text{new estimator}}\right)^{2}.$ It reflects how many times more MC samples are required to achieve the same accuracy as our new estimator. Section 5.3 shows that antithetic noise produces more diverse images than independent noise while preserving quality, and Appendix B.2 explores an image editing example.
5.1 Pixel-wise statistics
We begin by evaluating four pixel-level statistics. These include the (i) pixel value mean, (ii) perceived brightness, (iii) contrast, and (iv) the image centroid, with definitions in Appendix B.6. These statistics are actively used in diffusion workflows for diagnosing artifacts and assessing reliability, with applications in detecting signal leakage (Lin et al., 2024; Everaert et al., 2024), out-of-distribution detection (Le Bellier and Audebert, 2024), identifying artifact-prone regions (Kou et al., 2024), and improving the reliability of weather prediction (Li et al., 2024a).
Setup: We study both unconditional and conditional diffusion models; details are in Appendix B.3:
For unconditional diffusion models, we evaluate pre-trained models on CIFAR-10  (Krizhevsky et al., 2009), CelebA-HQ  (Xia et al., 2021), and LSUN-Church  (Yu et al., 2015). For each dataset, 1,600 image pairs are generated under both PN and RR noise sampling with 50 DDIM steps.
For conditional diffusion models, we evaluate Stable Diffusion 1.5 on 200 prompts from the Pick-a-Pic (Kirstain et al., 2023) and DrawBench (Saharia et al., 2022) with classifier-free-guidance (CFG) scale 3.5, and DiT  (Peebles and Xie, 2023) on 32 ImageNet classes with CFG scale 4.0. For each class or prompt, 100 PN and RR pairs are generated with 20 DDIM steps. We also study the effect of CFG scale on both models in the Appendix  B.4  (Ho and Salimans, 2022).
Implementation: To ensure fair comparisons under equal sample budget:
For unconditional diffusion models: MC uses 3,200 independent samples, AMC ($k=2$) uses 1,600 antithetic pairs, and AMC ($k=8$) uses 400 independent negative correlated batches. QMC employs a Sobol’ point set of size 64 with 50 independent randomizations. Due to the dimensionality limits, QMC is restricted to CIFAR-10.
For conditional diffusion models: for each class or prompt, MC uses 200 independent samples, AMC uses 100 antithetic pairs, and QMC employs 8 Sobol’ points with 25 randomizations.
CIs for MC and AMC are constructed as described in Section 4. For QMC, we use a student-$t$ interval to ensure reliable coverage  (L’Ecuyer et al., 2023), with details in Appendix B.6.